OpenStudy (anonymous):

How many three-letter codes can be formed using the letters A, B, C, and D? Repeated letters are allowed.

OpenStudy (anonymous):

4p3

OpenStudy (anonymous):

sorry 3^4 is ans

OpenStudy (anonymous):

u thr casmac!!!!!!!!

OpenStudy (anonymous):

yes i am!!!!!!

OpenStudy (anonymous):

but thats not the correct answer!!!!!!

jimthompson5910 (jim_thompson5910):

\[3^4=81\] different codes

OpenStudy (anonymous):

i put that as the answer and it told me i was wrong!!!!

OpenStudy (anonymous):

ho ans 4^3=64

jimthompson5910 (jim_thompson5910):

oops my bad, answer is \[4^3=64\], mixed up terms

jimthompson5910 (jim_thompson5910):

ie 4*4*4 = 64

OpenStudy (anonymous):

@casec jst check tis one !